OpenStudy (anonymous):

what is the inverse function of f(x)=x^5+x^3+x

OpenStudy (anonymous):

wait i am doing

OpenStudy (anonymous):

ok let g(x) = f^(-1) can you follow me?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

im going to substitute -1 to the original function for g

OpenStudy (anonymous):

when f has inverse function, it is symmetry at y=x right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

when i change all the y to x and vice verse, im stuck at x=y^5+y^3+1. i don't know what to do next because it seems whatever I do, it ends with y's being all over both sides

OpenStudy (anonymous):

in this problem it is really difficult for you to solve if you do like that

OpenStudy (anonymous):

is there a way that i can find the inverse, or will the inverse not be a function?

OpenStudy (anonymous):

did you learn derivative and integration?

OpenStudy (anonymous):

yes. would i have to do one of them

OpenStudy (anonymous):

ok let g is inverse of f then f(g(x))=x right?

OpenStudy (anonymous):

then take the derivative both side then you can get f'(g(x))g'(x)=1

OpenStudy (anonymous):

im not following that part. would i be using the g(-1) from above

OpenStudy (anonymous):

follow me?

OpenStudy (anonymous):

no just follow me

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait a second i will arrange some more time then i will tell you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hey is it the full question?? any other thing?

OpenStudy (anonymous):

i want to see full length of the question

OpenStudy (anonymous):

i think maybe thatis not the full length of the question

OpenStudy (anonymous):

the question was to just find the inverse of the function f(x)=x^5+x^3+x.

OpenStudy (anonymous):

never can't find just that inverse function because we can't arrage it in y after change x to y, we can just know what is g'(a) or what is g(a) , a is some constant

OpenStudy (anonymous):

ok. so i guess the inverse would not be a function since that being the only given info and y and x cant be rearrange anymore

OpenStudy (anonymous):

i also investigate more to find out what i don't know . keep in touch~

OpenStudy (anonymous):

HEY i have looked up my book. so the conculusion is cannot reduce it after change x to y . so x=y^5 + y^3 + y is the inversefunction of initial one